# Understanding and Mitigating Cross-lingual Privacy Leakage via Language-specific and Universal Privacy Neurons

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) (Achiam et al., 2023; Team et al., 2023; Liu et al., 2024) have significantly transformed the field of natural language processing (NLP). Benefiting from large-scale pretraining on multilingual corpora, these models demonstrate remarkable abilities in understanding and generating text across a wide range of languages (Huang et al., 2023; Zhao et al., 2024a). However, LLMs trained on a huge amount of internet data face critical privacy challenges, as they may memorize and unintentionally leak private information, particularly personally identifiable information (PII) (Huang et al., 2022; Li et al., 2023; Nakka et al., 2024). Currently, many studies aim to mitigate privacy or PII leakage in LLMs through techniques such as machine unlearning or privacy neuron-based interventions. The former seeks to erase memorized information by fine-tuning on small batches of data (Jang et al., 2022; Liu et al., 2025), while the latter reduces the likelihood of private information being elicited by directly modifying or suppressing relevant neurons (Wu et al., 2023). It is notable that most existing studies on PII assume that both the (sensitive) training data and user queries are in English (Lukas et al., 2023a; Kim et al., 2023a).
@@
To bridge this gap, we need to first understand why cross-lingual privacy leakage works (even under the current monolingual privacy-preserving techniques). To answer the question, we first extend an existing dataset to cover multiple languages and different types of PII. Based on the multilingual dataset, we analyze cross-lingual privacy leakage through the lens of mechanistic interpretability, which aims to elucidate the internal workings of LLMs. Specifically, we use Logit Lens (Nostalgebraist, 2020) to trace information flows within LLMs to pinpoint where cross-lingual privacy leakage occurs. Generally, our analysis reveals that LLMs process private information in the middle layers, which is largely shared across languages. The risk of privacy leakage peaks in the final layers, where the model transitions to language-specific generation.
@@ 2 Related Works
PII in LLMs. The potential for LLMs to memorize and leak PII has been a growing concern. Early studies, such as Carlini et al. (2021), demonstrated that models such as GPT-2 can reproduce sensitive data, including names and phone numbers, through extraction attacks. Nasr et al. (2023) further showed that even aligned models such as ChatGPT remain vulnerable, with their divergence attack increasing the emission of training data 150 times. To mitigate such vulnerabilities, Wu et al. (2023) developed DEPN, a framework to detect and edit privacy-related neurons in pre-trained models. By neutralizing the activations of neurons linked to sensitive data, DEPN reduces PII exposure while preserving model performance. Recent work by Lukas et al. (2023b) highlights the persistence of PII leakage, with novel attacks extracting up to 10 times more PII sequences than existing methods, even with differential privacy. Tools such as ProPILE (Kim et al., 2023b) offer practical ways for data subjects to assess PII leakage by formulating prompts based on their own PII. Broader surveys, such as Yao et al. (2024), provide an overview of LLM privacy challenges but do not specifically address cross-lingual privacy leakage, underscoring the need for further research on LLMs.
@@
Interpretability for Multilingual LLMs. Understanding the internal mechanisms of multilingual LLMs is crucial to analyzing their behavior across various linguistic contexts. Recent studies have used interpretability techniques to probe how multilingual large language models process and represent information. Clark et al. (2019) analyzed intermediate representations in transformer models to understand attention mechanisms, providing insights into information flow across layers.  Mueller et al. (2022) found that syntactic agreement in autoregressive multilingual models is encoded by overlapping cross-lingual neurons, indicating shared representational mechanisms.Tang et al. (2024) identified key language-specific neurons, while Kojima et al. (2024) demonstrated that controlling these neurons can manipulate the model’s output language, enhancing the understanding of cross-lingual knowledge transfer. However, these works primarily focus on knowledge probing or language understanding tasks, with limited attention to privacy-related issues, such as the leakage of PII in cross-lingual settings.
@@ 4.1 Experimental Setting
Implementation Details. After 10 epochs of fine-tuning on English texts from the MPII dataset Yang et al. (2024b), we construct parallel question-answer prompts across multiple languages to probe for cross-lingual privacy leakage (further details are provided in Appendix A.2). In addition, we evaluate the performance of the existing privacy-preserving method, DEPN, under cross-lingual settings to evaluate its effectiveness beyond the English context. DEPN is a framework for detecting and editing privacy neurons related to English-language data in pretrained language models, aiming to reduce privacy leakage risks in the post-processing stage without compromising model performance. Given that private information, such as job titles and email addresses, often consists of multiple tokens, we adopt the token-level Mean Reciprocal Rank (MRR) metric (Wu et al., 2023) to quantify how well the model memorizes and ranks target PII tokens. Further details are provided in Appendix A.3
@@ 5.2 Information Flow Perspective
We fine-tune the models for 10 epochs using only the English texts in MPII. After fine-tuning, we evaluate the cross-lingual privacy leakage by prompting it in multiple languages. To identify high-risk instances, we compute the MRR for each input prompt (the higher, the riskier). Based on this, we select the top 3% of samples with the highest MRR and categorize them into two groups: (1) instances that are high-risk when prompted in English, and (2) instances that are high-risk when prompted in non-English languages. We then compare the layer-wise MRR of these two groups across English and non-English prompt settings. This analysis allows us to trace how the model processes private information across layers and how such information transitions between languages.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) (Achiam et al., 2023; Team et al., 2023; Liu et al., 2024) have significantly transformed the field of natural language processing (NLP) Yang et al. (2024c). Benefiting from large-scale pretraining on multilingual corpora, these models demonstrate remarkable abilities in understanding and generating text across a wide range of languages (Huang et al., 2023; Zhao et al., 2024a). However, LLMs trained on a huge amount of internet data face critical privacy challenges, as they may memorize and unintentionally leak private information, particularly personally identifiable information (PII) (Huang et al., 2022; Li et al., 2023; Nakka et al., 2024). Currently, many studies aim to mitigate privacy or PII leakage in LLMs through techniques such as machine unlearning or privacy neuron-based interventions. The former seeks to erase memorized information by fine-tuning on small batches of data (Jang et al., 2022; Liu et al., 2025; Wang et al., 2024; Tao et al., 2024; Wang et al., 2023), while the latter reduces the likelihood of private information being elicited by directly modifying or suppressing relevant neurons (Wu et al., 2023). It is notable that most existing studies on PII assume that both the (sensitive) training data and user queries are in English (Lukas et al., 2023a; Kim et al., 2023a).
@@
To bridge this gap, we need to first understand why cross-lingual privacy leakage works (even under the current monolingual privacy-preserving techniques). To answer the question, we first extend an existing dataset to cover multiple languages and different types of PII. Based on the multilingual dataset, we analyze cross-lingual privacy leakage through the lens of mechanistic interpretability, which aims to elucidate the internal workings of LLMs. Specifically, we use Logit Lens (Nostalgebraist, 2020; Zhang et al., 2024) to trace information flows within LLMs to pinpoint where cross-lingual privacy leakage occurs. Generally, our analysis reveals that LLMs process private information in the middle layers, which is largely shared across languages. The risk of privacy leakage peaks in the final layers, where the model transitions to language-specific generation.
@@ 2 Related Works
PII in LLMs. The potential for LLMs to memorize and leak PII has been a growing concern. Early studies, such as Carlini et al. (2021), demonstrated that models such as GPT-2 can reproduce sensitive data, including names and phone numbers, through extraction attacks. Nasr et al. (2023) further showed that even aligned models such as ChatGPT remain vulnerable, with their divergence attack increasing the emission of training data 150 times. To mitigate such vulnerabilities, Wu et al. (2023) developed DEPN, a framework to detect and edit privacy-related neurons in pre-trained models. By neutralizing the activations of neurons linked to sensitive data, DEPN reduces PII exposure while preserving model performance. Recent work by Lukas et al. (2023b) highlights the persistence of PII leakage, with novel attacks extracting up to 10 times more PII sequences than existing methods, even with differential privacy Huang et al. (2024); Xiao et al. (2023); Xiang et al. (2023); Wang et al. (2019); Wang and Xu (2019); Zhang et al. (2025a). Tools such as ProPILE (Kim et al., 2023b) offer practical ways for data subjects to assess PII leakage by formulating prompts based on their own PII. Broader surveys, such as Yao et al. (2024); Hu et al. (2024a), provide an overview of LLM privacy challenges but do not specifically address cross-lingual privacy leakage, underscoring the need for further research on LLMs.
@@
Interpretability for Multilingual LLMs. Understanding the internal mechanisms of multilingual LLMs is crucial to analyzing their behavior across various linguistic contexts. Recent studies have used interpretability techniques to probe how multilingual large language models process and represent information Hong et al. (2024); Yao et al. (2025); Zhang et al. (2025b); Hu et al. (2024c). Hu et al. (2025, 2024b, 2024d); Lai et al. (2023); Hu et al. (2023) studied the stability of the explanation given by the attention mechanism. Clark et al. (2019) analyzed intermediate representations in transformer models to understand attention mechanisms, providing insights into information flow across layers.  Mueller et al. (2022) found that syntactic agreement in autoregressive multilingual models is encoded by overlapping cross-lingual neurons, indicating shared representational mechanisms.Tang et al. (2024) identified key language-specific neurons, while Kojima et al. (2024) demonstrated that controlling these neurons can manipulate the model’s output language, enhancing the understanding of cross-lingual knowledge transfer. However, these works primarily focus on knowledge probing or language understanding tasks, with limited attention to privacy-related issues, such as the leakage of PII in cross-lingual settings.
@@ 4.1 Experimental Setting
Implementation Details. After 10 epochs of fine-tuning on English texts from the MPII dataset, we construct parallel question-answer prompts across multiple languages to probe for cross-lingual privacy leakage (further details are provided in Appendix A.2). In addition, we evaluate the performance of the existing privacy-preserving method, DEPN, under cross-lingual settings to evaluate its effectiveness beyond the English context. DEPN is a framework for detecting and editing privacy neurons related to English-language data in pretrained language models, aiming to reduce privacy leakage risks in the post-processing stage without compromising model performance. Given that private information, such as job titles and email addresses, often consists of multiple tokens, we adopt the token-level Mean Reciprocal Rank (MRR) metric (Wu et al., 2023) to quantify how well the model memorizes and ranks target PII tokens. Further details are provided in Appendix A.3
@@ 5.2 Information Flow Perspective
We fine-tune the models for 10 epochs using only the English texts in MPII. After fine-tuning Yang et al. (2024b), we evaluate the cross-lingual privacy leakage by prompting it in multiple languages. To identify high-risk instances, we compute the MRR for each input prompt (the higher, the riskier). Based on this, we select the top 3% of samples with the highest MRR and categorize them into two groups: (1) instances that are high-risk when prompted in English, and (2) instances that are high-risk when prompted in non-English languages. We then compare the layer-wise MRR of these two groups across English and non-English prompt settings. This analysis allows us to trace how the model processes private information across layers and how such information transitions between languages.